# Structure-preserving space discretization of differential and nonlocal constitutive relations for port-Hamiltonian systems

### v1

@@ 5.1 Nanorod equation
Let us now study the numerical solution of the nanorod equation obtained after discretization. We will consider the domain $\Omega=[0,1]$, the Young’s modulus $E=1$, the mass density $\rho=10$, and as initial conditions, we choose $v_{0}(x)=\exp(-80(x-0.3)^{2})$ and $\sigma_{0}(x)=0$. We will then study the effect of the parameter $\ell$ over 00, $0.01$, and $0.05$. Finally, the implicit midpoint rule is chosen as the time scheme. The simulation parameters are: final time $T_{f}=10$, time step $\textup{d}t=0.1$ and number of discretization points $N=100.$ Figure 2 shows of the evolution of the Hamiltonian for the different values of $\ell$, along with respective relative energy errors $|H_{\text{Rob}}^{d}(t)-H_{\text{Rob}}^{d}(0)|/H_{\text{Rob}}^{d}(0)$.
In each setting, the model is conservative hence the total mechanical energy should remain constant, however, it is clear that increasing the parameter $\ell$ decreases the accuracy of the method. Such behaviour might be caused by the degradation of the condition number of the matrix $\bm{M}+\ell^{2}\bm{K}+\ell\bm{BB}^{\top}$ as $\ell$ is increased (see C.1). Additionaly, on the top row, the green and red lines isolate the energy contribution of the velocity energy port $\ell\,\overline{v}\bm{BB}^{\top}\overline{v}$ and stress energy port $\frac{\ell}{2E}\,\overline{\sigma}\bm{BB}^{\top}\overline{\sigma}$ respectively. These show that the amount of energy stored at the boundaries becomes significant as $\ell$ increases and should not be omitted when considering Robin boundary conditions.
@@ 5.2 Implicit Euler-Bernoulli beam
Figure 3: Phase velocity for $r=5~{}$cm (left) and $r=2.5~{}$cm (right), and a mesh size parameter $\textup{d}x=5.0\times 10^{-4}$ (implicit Euler-Bernoulli beam).
@@
In Figure 6, the preservation of the underlying Dirac and Lagrange structures are enlightened: the variations of Hamiltonian are close to machine precision. One may furthermore appreciate that the implicit parameter, while improving the quality of the simulation for high frequencies, the requirement of CPU time is not significantly different. This latter is mainly due to the better condition number of the matrix $\rho h\bm{M}+\frac{\rho h^{3}}{12}\bm{K}$ on the left-hand side of (43), when $h$ belongs to an appropriate range, because this results in a larger timestep, which counterbalances the time spent for the resolution of the linear system involving $\bm{K}$. Indeed, it has already been explained in [26] that the implicit Euler-Bernoulli beam (the shear model), is well-suited for medium-sized beams, while Timoshenko’s model is better suited for thick beams and (explicit) Euler-Bernoulli’s for thin beams.
@@
5.3 Incompressible Navier-Stokes equation
@@ 5.3.1 Dipole collision
Moreover, boundary conditions are chosen as no-slip and impermeable. As a consequence, the initial condition $\psi_{0}$ is chosen to be the solution of the Poisson equation $-\Delta\psi_{0}=\omega_{0}$ with homogeneous Dirichlet boundary conditions to enforce the impermeability. As highlighted in [23], with such initial conditions, the tangential velocity corresponding to $\omega_{0}$ at the boundaries is very small, hence $\psi_{0}$ satisfies the no-slip boundary condition as well.
@@
Figure 10 presents the vorticity profile at the right boundary over the times 0.4 s, 0.6 s and 1.0 s. This shows how vorticity is being generated by the Lagrange multiplier $u_{D}^{3}=\omega_{|\partial\Omega}$ to enforce the no-slip boundary condition. Moreover, figure 10 presents a contour plot of the vorticty on the subdomain $[0.4,1]\times[0,0.6]$ at time $t=1$s showing how the upper part of the dipole interacts with the boundary after initial collision.
@@
Figure 10: Vorticity profile at the right boundary $x=1$ and along $y\in[-0.6,0]$ for the times $0.4$s, $0.6$s and $1.0$s (left). Vorticity contour plot at $t=1$s in the subdomain $[0.4,1]\times[0,0.6]$ (right).
@@ A.1 Proof of Lemma 2
Moreover, from $P^{\top}S-S^{\top}P=0$, it comes: $\text{Ran}\begin{bmatrix}P\\ S\end{bmatrix}\subset\ker(\begin{bmatrix}S^{\top}&-P^{\top}\end{bmatrix}).$
Furthermore, we have from the size of $P$ and $S$: $\text{dim}(\ker(\begin{bmatrix}S^{\top}&-P^{\top}\end{bmatrix})\leq n+n_{L}$, and the rank condition gives $\text{dim}(\text{Ran}\begin{bmatrix}P\\ S\end{bmatrix})=n+n_{L}$ Hence,
@@ A.2 Proof of Lemma 3
Let us compute $\frac{\textup{d}}{\textup{d}t}H(\lambda,\tilde{u})$:
$\frac{\textup{d}}{\textup{d}t}H(\lambda,\tilde{u})$ $=\frac{\textup{d}}{\textup{d}t}\left(\frac{1}{2}\begin{pmatrix}\lambda\\ \tilde{u}_{L}\end{pmatrix}^{\top}P^{\top}S\begin{pmatrix}\lambda\\ \tilde{u}_{L}\end{pmatrix}\right)$
@@
00 0 $=\begin{pmatrix}e\\ e_{R}\\ u_{D}\end{pmatrix}^{\top}J\begin{pmatrix}e\\ e_{R}\\ u_{D}\end{pmatrix}$ $(J=-J^{\top})$
@@ A.3 Proof of Lemma 7
$\begin{array}[]{rcl}\sigma(a)+n_{a}\,\ell\partial_{x}\sigma(a)=0,&&\sigma(b)+n_{b}\,\ell\partial_{x}\sigma(b)=0,\\ \lambda(a)+n_{a}\,\ell\partial_{x}\lambda(a)=0,&&\lambda(b)+n_{b}\,\ell\partial_{x}\lambda(b)=0.\end{array}$
@@ C.1 Condition number
Figure 11: Condition number of matrix $\bm{M}+\ell^{2}\,\bm{K}+\ell\,\bm{B}\bm{B}^{\top}$ for number of discretization points $N=$100, 500 and 1000.
Appendix D Incompressible Navier-Stokes equation
D.1 Time discretization

### v2

@@ 5.1 Nanorod equation
Let us now study the numerical solution of the nanorod equation obtained after discretization. We will consider the domain $\Omega=[0,1]$, the Young’s modulus $E=1$, the mass density $\rho=10$, and as initial conditions, we choose $v_{0}(x)=\exp(-80(x-0.3)^{2})$ and $\sigma_{0}(x)=0$. We will then study the effect of the parameter $\ell$ over 0, $0.01$, and $0.05$. Finally, the implicit midpoint rule is chosen as the time scheme. The simulation parameters are: final time $T_{f}=10$, time step $\textup{d}t=0.1$ and number of discretization points $N=100.$ Figure 2 shows of the evolution of the Hamiltonian for the different values of $\ell$, along with respective relative energy errors $|H_{\text{Rob}}^{d}(t)-H_{\text{Rob}}^{d}(0)|/H_{\text{Rob}}^{d}(0)$.
In each setting, the model is conservative hence the total mechanical energy should remain constant, however, it is clear that increasing the parameter $\ell$ decreases the accuracy of the method. Such behaviour might be caused by the degradation of the condition number of the matrix $\bm{M}+\ell^{2}\bm{K}+\ell\bm{BB}^{\top}$ as $\ell$ is increased (see C.1). Additional ly, on the top row, the green and red lines isolate the energy contribution of the velocity energy port $\ell\,\overline{v}\bm{BB}^{\top}\overline{v}$ and stress energy port $\frac{\ell}{2E}\,\overline{\sigma}\bm{BB}^{\top}\overline{\sigma}$ respectively. These show that the amount of energy stored at the boundaries becomes significant as $\ell$ increases and should not be omitted when considering Robin boundary conditions.
@@ 5.2 Implicit Euler-Bernoulli beam
Figure 3: Phase velocity for r=5r=5~cm (left) and r=2.5r=2.5~cm (right), and a mesh size parameter $\textup{d}x=5.0\times 10^{-4}$ (implicit Euler-Bernoulli beam).
@@
In Figure 6, the preservation of the underlying Dirac and Lagrange structures are enlightened: the variations of Hamiltonian are close to machine precision. One may furthermore appreciate that the implicit parameter, while improving the quality of the simulation for high frequencies, the requirement of CPU time is not significantly different. This latter is mainly due to the better condition number of the matrix $\rho h\bm{M}+\frac{\rho h^{3}}{12}\bm{K}$ on the left-hand side of (42), when $h$ belongs to an appropriate range, because this results in a larger timestep, which counterbalances the time spent for the resolution of the linear system involving $\bm{K}$. Indeed, it has already been explained in [26] that the implicit Euler-Bernoulli beam (the shear model), is well-suited for beams of moderate thickness, while Timoshenko’s model is better suited for thick beams and (explicit) Euler-Bernoulli’s for thin beams.
@@
5.2.1 Convergence analysis
Let us study the convergence of the previously discretized model. Here, we will look at the phase velocity of the first 50 eigenvalues of the implicit Euler-Bernoulli beam. For a mode $i$ we denote by $c_{p,i}$ the corresponding phase velocity and $k_{i}$ the corresponding wavenumber. As such, we study the convergence of the relative error
$\frac{1}{50}\sum_{i=1}^{50}\frac{|c_{p,i}-c_{p}^{th}(k_{i})|}{c_{p}^{th}(k_{i})}$
with respect to the mesh step size $\textup{d}x$, with $c_{p}^{th}(k)=k\sqrt{D}/\sqrt{\rho h(1+(h^{2}/12)k^{2}}$. The model parameter are chosen as : $D=5\times 10^{5},\rho=8\times 10^{3},h=6.28\times 10^{-2}$. Figure 7 shows the convergence of the phase velocities with respect to the step size $\textup{d}x,$ in particular we obtain a convergence order of 1.106.
Figure 7: Convergence of the phase velocities for the implicit Euler-Bernoulli model
5.3 Incompressible Navier-Stokes equations
@@ 5.3.1 Dipole collision
Moreover, boundary conditions are chosen as no-slip and impermeable. As a consequence, the initial condition $\psi_{0}$ is chosen to be the solution of the Poisson equation $\Delta\psi_{0}=\omega_{0}$ with homogeneous Dirichlet boundary conditions to enforce the impermeability. As highlighted in [23], with such initial conditions, the tangential velocity corresponding to $\omega_{0}$ at the boundaries is very small, hence $\psi_{0}$ satisfies the no-slip boundary condition as well.
@@
Figure 11 presents the vorticity profile at the right boundary over the times 0.4 s, 0.6 s and 1.0 s. This shows how vorticity is being generated by the Lagrange multiplier $u_{D}^{3}=\omega_{|\partial\Omega}$ to enforce the no-slip boundary condition. Moreover, figure 11 presents a contour plot of the vorticty on the subdomain $0.4,1]\times[0,0.6]$ at time $t=1$s showing how the upper part of the dipole interacts with the boundary after initial collision.
@@
Figure 11: Vorticity profile at the right boundary $x=1$ and along $y\in[-0.6,0]$ for the times $0.4$s, $0.6$s and $1.0$s (left). Vorticity contour plot at $t=1$s in the subdomain $0.4,1]\times[0,0.6]$ (right).
5.3.2 Convergence analysis
Let us study the convergence of the incompressible Navier-Stokes equations when refining the mesh. We first deal with Taylor-Green vortices, where the solution is mainly driven by the dissipative term and an analytical solution is known ; then we address the normal dipole collision, where the convective term is no more negligible, and the finest mesh is used as a reference since no analytical solution is known.
Taylor-Green vortices
First, let us look at an analytical solution, namely the Taylor-Green vortices case [24]:
$\psi_{ref}(t,x,y)$ $=\frac{1}{\pi}\sin(\pi x)\sin(\pi y)\exp(-\frac{2\pi^{2}t}{Re}),$
$\omega_{ref}(t,x,y)$ $=-\Delta\psi_{ref}(t,x,y)=2\pi^{2}\psi_{ref}(t,x,y),$
for $(x,y)\in[0,1]^{2},t\in\mathbb{R}^{+},$ with $Re=\rho_{0}/\mu,$ the Reynolds number ; in our case $\rho_{0}=1,$ and $\mu=1/100$. Moreover, the boundary conditions are homogeneous Dirichlet for both $\omega$ and $\psi$. The mesh is taken as a regular grid of simplexes (triangles) composed of $K$ rows and $K$ columns with $K\in\{5,7,9,11,13,15,17,19,21,23,25\}$. The time step is taken as $\textup{d}t=1/1000s$ and final time is $t=1s.$
Figure 12 shows the $H^{1}$ error $\omega-\omega_{ref}||_{H^{1}}$ at time $t=1s$ and $\psi-\psi_{ref}||_{H^{1}}$ at time $t=0.9995s$ (since $\psi$ is evaluated on half time steps) respectively, plotted against the average element size of the mesh (triangle circumradius).
In such a case, we recover a convergence order of $2.565$ for $\omega$ and $4.760$ for $\psi$ with respect to the average element sizes.
Note however that with such a choice of variables (vorticity/stream function) yields a solution mainly driven by the dissipative term and not the convective term. Indeed, the convective term is identically zero:
$\operatorname{div}(\omega_{ref}\,\textbf{{grad}}^{\perp}(\psi_{ref}))=\textbf{{grad}}(\omega_{ref})\cdot\textbf{{grad}}^{\perp}(\psi_{ref})=2\pi^{2}\,\textbf{{grad}}(\psi_{ref})\cdot\textbf{{grad}}^{\perp}(\psi_{ref})=0.$
Figure 12: Vorticity and stream function $H^{1}$ error for the Taylor-Green vortices case.
Dipole collision
Let us now study the convergence of the normal dipole collision case when refining the mesh. Here, the convective term is not negligible. To do so, we ran the simulation with $\rho_{0}=1,\mu=625,\textup{d}t=1/600$, a final time $t=1.5s$ and for 9 different meshes with a mesh size parameter $\ell$ ranging from $0.04$ to $0.12$ and an average triangle size ranging from $0.0055$ to $0.016$. The finest mesh ($\ell=0.04$) is used as a reference. The error is then defined as the average of the relative $H^{1}$ errors for 180 evenly distributed timesteps between 0 and $1.5s$:
[…]
We obtain a convergence order of 2.234 and 3.910 for $\omega$ and $\psi$ respectively. Figure 13 shows the error for both $\psi$ and $\omega$ plotted against the average element size.
Figure 13: Vorticity and stream function $H^{1}$ relative error for the dipole collision case.
@@ A.1 Proof of Lemma 2
[…]
Furthermore, we have from the size of $P$ and $S$: $\text{dim}(\ker(\begin{bmatrix}S^{\top}&-P^{\top}\end{bmatrix})\leq n+n_{L}$, and the rank condition gives $\text{dim}(\text{Ran}\begin{bmatrix}P\\ S\end{bmatrix})=n+n_{L}$ . Hence,
@@ A.2 Proof of Lemma 3
[…]
@@
$0$ $=\begin{pmatrix}e\\ e_{R}\\ u_{D}\end{pmatrix}^{\top}J\begin{pmatrix}e\\ e_{R}\\ u_{D}\end{pmatrix}$ $(J=-J^{\top})$
@@ A.3 Proof of Lemma 8
$\begin{array}[]{rcl}\sigma(a)+n_{a}\,\ell\partial_{x}\sigma(a)=0,&\qquad&\sigma(b)+n_{b}\,\ell\partial_{x}\sigma(b)=0,\\ \lambda(a)+n_{a}\,\ell\partial_{x}\lambda(a)=0,&\qquad&\lambda(b)+n_{b}\,\ell\partial_{x}\lambda(b)=0.\end{array}$
@@ C.1 Condition number
Figure 14: Condition number of matrix $\bm{M}+\ell^{2}\,\bm{K}+\ell\,\bm{B}\bm{B}^{\top}$ for number of discretization points N=N=100, 500 and 1000.
Appendix D Implicit Euler-Bernoulli beam
Let us derive the model following [52, Chapter 2], where thin plate models are tackled. We neglect the $y$ component to model the beam case of this work.
Let us denote by $U(x,z,t)$ the displacement at a time $t$ of the particle that occupies the position $(x,z)$ when the beam is at equilibrium. Furthermore, we denote $u(x,t)=U(x,0,t)$, the displacement of the middle line.
Assume that the strain is proportional to the displacement (i.e., a linear strain displacement relation), and that the linear filaments of the beam initially perpendicular to the middle line remain straight and perpendicular to the deformed middle line, and undeformable: neither contraction nor extension (transverse shear effects are neglected).
The second assumption imposes a nonlinear relationship between $U(x,z,t)$ and $u(x,t)$. If this relation is linearized, we obtain: $U=u-z\partial_{x}u_{z}$, where $u_{z}$ denotes the $z$ component of the vector field $u$.
Denoting $w:=u_{z}$, the vertical displacement, the strain energy $\mathcal{P}$ of the beam is defined by:
where L:=baL:=b-a is the length of the beam and $D:=E\frac{h^{3}}{12(1-\nu^{2})}$ is the modulus of flexural rigidity, with $E$ the Young’s modulus, $\nu$ the Poisson’s ratio ($0<\nu<\frac{1}{2}$ in physical situations), and $h$ the cross-sectional area ($h:=\pi r^{2}$ for cylindric beam of radius $r$).
Note that to include a tension $T_{0}$, as done in [9, Eq. (10)], and model pre-stressed beams, it requires adding more potential energy terms, following [26, Eq. (2)]. As this is not the purpose of this work, we restrict ourselves to the simpler models without tension.
The kinetic energy reads:
Principle of virtual work
$\mathcal{L}=\int_{0}^{T}\mathcal{K}+\mathcal{W}-\mathcal{P}\,\textup{d}t,$
More precisely, let $\eta=\delta w$ be any admissible variation of $w$ in $\mathcal{C}([0,T]\times[a,b])$ with compact support, then:
$\delta\mathcal{P}=D\int_{a}^{b}\frac{\partial^{4}}{\partial x^{4}}w\,\eta\,\textup{d}x,$
$\int_{0}^{T}\delta\mathcal{K}\,\textup{d}t=\rho h\int_{0}^{T}\int_{a}^{b}\left(-\frac{\partial^{2}}{\partial t^{2}}w+\frac{h^{2}}{12}\frac{\partial^{4}}{\partial x^{2}\partial t^{2}}w\right)\eta\,\textup{d}x\,\textup{d}t,$
and
$\delta\mathcal{W}=\int_{a}^{b}f\,\eta\,\textup{d}x,$
where $f$ is the external force.
Since $\eta$ is arbitrary, the equation of motion for the Euler-Bernoulli beam is obtained by setting $\delta\mathcal{L}=\int_{0}^{T}\delta\mathcal{K}+\delta\mathcal{W}-\delta\mathcal{P}\,\textup{d}t=0$, which then reads:
$\rho h\left(1-\frac{h^{2}}{12}\frac{\partial^{2}}{\partial x^{2}}\right)\frac{\partial^{2}}{\partial t^{2}}w+D\frac{\partial^{4}}{\partial x^{4}}w=f.$ (54)
Appendix E Incompressible Navier-Stokes equation
E.1 Time discretization